OpenStudy (anonymous):

i am having problems with integration of rational functions by partial fractions. the question is z+1/z^2(z-1)

OpenStudy (freckles):

\[\frac{z+1}{z^2(z-1)}=\frac{A}{z}+\frac{B}{z^2}+\frac{C}{z-1}\] find A,B, and C such that equation is true

OpenStudy (anonymous):

finding A, B and C is what i need help with

OpenStudy (freckles):

do you know how to combine fractions?

OpenStudy (freckles):

look at the first fraction A/z what do we need on bottom so that eventually we will have the same denominators well it is lacking another z and a (z-1) so you will multiply top and bottom by z(z-1) --- look at the second fraction B/z^2 what do we need on bottom? well it is lacking a (z-1) so we will multiply top and bottom by (z-1) --- look at C/(z-1) what do we need on bottom? well it is lacking a z^2 so we will multiply top and bottom by z^2

OpenStudy (freckles):

\[\frac{z+1}{z^2(z-1)}=\frac{A}{z}+\frac{B}{z^2}+\frac{C}{z-1} \\ \frac{z+1}{z^2(z-1)}=\frac{A z(z-1)+B(z-1)+Cz^2}{z^2(z-1)}\]

OpenStudy (freckles):

bottoms are the same just compare the tops

OpenStudy (freckles):

you have to satisfy the equation \[z+1=A(z^2-z)+B(z-1)+Cz^2\]

OpenStudy (freckles):

collect your like terms together

OpenStudy (anonymous):

You need to get the constants A,B,C shown in the next ecuation. \[z+1/z ^{2}(z-1)= (A/z-1) + (B/z) + (C/z ^{2})\] Multiplying each side by z^2(z-1) \[z+1=z ^{2}(z-1)(A/z-1)+z ^{2}(z-1)(B/z)+z ^{2}(z-1)(C/z ^{2})\] After that you can give z any value that could hel you to solve the ecuation and get A,B,C. Example: z=1, z=0. Both of them make the ecuation easier to solve and give you constant values. After that evaluate with the constants and solve the 3 integers. You should obtamin solutions with logarithms.

OpenStudy (freckles):

\[z+1=(A+C)z^2+(-A+B)z+(-B)\]

OpenStudy (freckles):

notice you have 0 z^2 's on left hand side notice you have (A+C) z^2 's on the right hand side so you need A+C to be 0 So you have A+C=0 -- notice you have 1 z 's on left side notice you have (-A+B) z's on right side so you need 1=-A+B --- notice you have constant 1 on left side notice you have constant -B on right side so you need -B=1

OpenStudy (freckles):

last equation solve it first then use it to solve second equation then use that result and solve the first equation

OpenStudy (anonymous):

your wrong because the answer is -2/z + -1/z^2 + 2/z-1

OpenStudy (freckles):

who are you talking to?

OpenStudy (freckles):

you have to solve the equations...

OpenStudy (freckles):

you can do this if -B=1 then B=?

OpenStudy (freckles):

if -B=1 then B=-1 now go to the second equation 1=-A+B you have B=-1 solve for A

OpenStudy (anonymous):

you and no, from my teacher, in class told me the equation is z+1=Az(z-1)+B(z-1)+ Cz^2, i have C=2, A=-1 but i cant find B

OpenStudy (freckles):

Your A value in incorrect

OpenStudy (anonymous):

i mean C=2, A=-1 and i dont know B

OpenStudy (freckles):

anyways i don't know what you are calling wrong but I have not said anything incorrect I gave you the three equations to solve all I did was compare both sides after combining like terms and got the following three equations: A+C=0 -A+B=1 -B=1

OpenStudy (freckles):

The last equation is the easiest equation to solve if -B=1 then B=-1 then use the equation equation -A+B=1 replace B with -1 -A+(-1)=1 solve for A -A=2 A=-2 then go to the first equation A+C=0 -2+C=0 which gives C=2

OpenStudy (freckles):

now you just go back to what were trying to split up as before \[\frac{z+1}{z^2(z-1)}=\frac{A}{z}+\frac{B}{z^2}+\frac{C}{z-1}\]

OpenStudy (anonymous):

yeah i have that, so there for \[you would have z+1=Az(z-1)+B(z-1)+cz ^{2}\]

OpenStudy (freckles):

yes same equation I have way above somewhere

OpenStudy (freckles):

you do know z(z-1) is the same as z^2-z so that is not wrong

OpenStudy (freckles):

still don't know what you refer to as wrong

OpenStudy (anonymous):

because you get a common denominator

OpenStudy (freckles):

saying your wrong helps me in no way to help you you could say if you don't understand something I did

OpenStudy (freckles):

but I would like to know what it is

OpenStudy (anonymous):

i quess my teacher does it a little differently than you, lets start over, i already have \[A/z+B/z ^{2}+C/(z-1)\]

OpenStudy (freckles):

ok and you said you were able to get \[z+1=A(z^2-z)+B(z-1)+Cz^2\]

OpenStudy (freckles):

now I said something about combining any like terms ... \[z+a=(A+C)z^2+(-A+B)z+(-B)\] do you understand how I rearranged things here?

OpenStudy (freckles):

that lowercase a is suppose to be 1

OpenStudy (freckles):

\[z+1 =(A+C)z^2+(-A+B)z+(-B)\]

OpenStudy (anonymous):

yes i see what you did there, now in order to find c you, you put like a 1 in for z and solve,. and the same with the others, just with low numbers like -1 or 0

OpenStudy (freckles):

no

OpenStudy (freckles):

i didn't do that

OpenStudy (freckles):

i compared both sides of the equation

OpenStudy (freckles):

as i said before there are 0 z^2's on the right but (A+C) z^2's on the left so 0=A+C and there is 1 z on the right and (-A+B) z on the left so 1=-A+B and there there is 1 on the right and -B on the left so 1=-B

OpenStudy (freckles):

but if you want to do it that other way we can try normally i go for solving the system of equations

OpenStudy (anonymous):

i just learned this last week so i am a little lost on your method

OpenStudy (freckles):

\[z+1=Az(z-1)+B(z-1)+Cz^2 \\ \\ z=1 \text{ gives } 1+1=A(1)(1-1)+B(1-1)+C(1)^2 \\ 2=A(0)+B0)+C \\ C=2 \text{ which is the same value I got with the previous method } \\ z=0 \text{ gives } 0+1=A(0)(0-1)+B(0-1)+2(0)^2 \\ \text{ so here you get } 1=B(-1) \\ \text{ or } -1 =B \text{ which is also what I have by previous method } \\ \text{ do you understand this so far }\]

OpenStudy (freckles):

We only have to find A now

OpenStudy (freckles):

we have C=2 and B=-1 so we have so far the following equation: \[z+1=Az(z-1)-1(z-1)+2z^2 \]

OpenStudy (freckles):

use some value that is not 0 or 1 and solve for A

OpenStudy (anonymous):

ok, i have been using -2, -1, -+.5 and getting bad answers

OpenStudy (freckles):

how did you plus or mins 1/2

OpenStudy (anonymous):

those are the values i plugged in for z

OpenStudy (freckles):

i'm going to replace z with -1... \[z+1=Az(z-1)-1(z-1)+2z^2 \\ -1+1=A(-1)(-1-1)-1(-1-1)+2(-1)^2 \\ 0=A(-1)(-2)-1(-2)+2(1) \\ 0=2A+2+2 \\ \text{ subtract 4 on both sides } -4=2A \text{ divide by 2 on both sides } -2=A\] which is the result i got by previous method

OpenStudy (anonymous):

i must of done my rhythmic wrong thank you for you patience, calc 2 is frustrating me

OpenStudy (freckles):

just a pointer you may not want to refer to someone 's work as wrong without even knowing anything about that method it just comes off as negative and kinda rude

OpenStudy (anonymous):

im sorry

OpenStudy (freckles):

it's okay i was just a tad bit annoyed

OpenStudy (freckles):

anyways try to go back through the problem or go back to my work and please let me know if I'm can clarify anything I did you can look at both methods or not it is up to you

OpenStudy (freckles):

I did start with z=1 because it made two terms 0 then I when to z=0 because it made one term 0 then I could have chose any z for the last one but I thought z=-1 would make things easiest

OpenStudy (anonymous):

ok, well thank you

OpenStudy (freckles):

np